# Associative Knowledge Graphs for Efficient Sequence Storage and Retrieval

### v1

@@ 7 Conclusions
We found that the memory capacity is determined by the size of the knowledge graph and the density of its synaptic connections. Our theoretical analysis reveals that memory capacity grows quadratically with the number of neurons used to build the knowledge graph. While smaller graphs may not yield significant sequence capacity, larger graphs can store a considerable number of sequences, allowing each neuron to represent multiple sequences.
@@
We conducted memory tests using both randomly generated synthetic data and real-world datasets, such as a collection of sequences of words featuring different sentences. These tests confirmed the validity of our findings.
@@
Our future work related to this topic will focus on two key aspects. First, we aim to compare the proposed structural associative memory with contemporary associative memory solutions, such as LSTM or NuPick. A comparative analysis of these two solutions will be the subject of our upcoming article. Second, we plan to enhance the presented node ordering concept by incorporating novel elements like using prime numbers to represent sequence order and micro-columns to increase memory capacity.

### v2

@@ 4 Discussion and conclusion
We found that the memory capacity is determined by the size of the knowledge graph and the density of its synaptic connections. Our theoretical analysis reveals that memory capacity grows quadratically with the number of neurons used to build the knowledge graph. While smaller graphs may not yield significant sequence capacity, larger graphs can store a considerable number of sequences, allowing each node to represent multiple sequences.
If the stored sequences have many elements in common, the result of resolving them depends on the context used. If the context is common to several stored sequences, then the algorithm will return all elements contained in several sequences. Such an extended sequence cannot be sorted properly. Such a result is considered incorrect. In our tests, we only considered it a success if the entire sequence is returned without error.
The method of MicroRNA (miRNA) coding described in Example 5 minimizes this behavior of the algorithm. In all MicroRNA tests, sequences were selected from the database at random.
@@
We conducted memory tests using both randomly generated synthetic data and real-world datasets, such as a collection of sequences of words featuring different sentences. Additionally, we tested an example using a genetic database. These tests confirmed the validity of our findings.
@@
  • The authors provide an open-source implementation of SSAKGs via the SSAKG Python package, along with Jupiter notebooks and tutorials.
  • This accessibility promotes adoption and further exploration of the methodology in both academic and clinical settings.
  • Biomedical researchers can adapt these techniques for anomaly detection, disease progression modeling, or personalized medicine.
Compared to traditional Hopfield networks, which rely on a fixed number of neurons to store memories, this method offers a more efficient solution. By utilizing sparse knowledge graphs and contextual connections, it can store and recall a greater number of scenes without sacrificing performance.
The method’s limitations stem from the theoretical constraints expressed in equations 6 and 7. The memory capacity, which is largely influenced by the density of the knowledge graph and the length of memorized sequences, is significantly reduced as the sequences grow longer. This limitation is directly linked to the parameter $\xi$ in equation 5, which increases with increasing sequence length. This approach has broad applications, from computational neuroscience to bioinformatics, offering scalable solutions for sequence-based memory tasks.
Our future work related to this topic will focus on two key aspects. First, we aim to compare the proposed structural associative memory with contemporary associative memory solutions, such as Modern Hopfield Networks (MHN), LSTM, and NuPIC. A comparative analysis of these two solutions will be the subject of our upcoming article. Second, we plan to enhance the presented node ordering concept by incorporating novel elements like using prime numbers to represent sequence order and micro-columns to increase memory capacity.
Credit authorship contribution statement
Przemysław Stokłosa: Software design, writing, Methodology, Conceptualization, Writing – review & editing, Project administration, Data acquisition, Resources. Janusz A. Starzyk: Methodology, Conceptualization, Software writing, Writing – review & editing, Data acquisition. Paweł Raif: Writing – review & editing, Resources, Project administration, Software tests, Funding acquisition. Adrian Horzyk: Writing – review & editing, Resources. Marcin Kowalik: Writing – review & editing, Data acquisition, Resources.
Supplementary materials
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.